At 12.25 degree C a brass sleeve has an inside dia
Answer #1

a)

the original length of the sleeve is

L_o=D.\pi=2.196cm*\pi

the final length

L=D.\pi=2.199cm*\pi

\Delta L=2.199cm*\pi-2.196cm*\pi=0.003\pi

\Delta L=\alpha (L_o)\Delta T

clearing delta T

\Delta T=\frac{\Delta L}{\alpha (L_o)}

\Delta T=\frac{0.003\pi cm}{1.9\times 10^{-5}K^{-1}(2.196\pi cm)}=71.9ºK

T_f=T_o+71.9ºK\Rightarrow T_f=84.15^0C

anserw a= 84.15ºc

/************************************************************

The shaft is an expansion of area. it could be calculated by the formula

\Delta A=2\alpha (A_o)\Delta T

A_o=\frac{\pi}{4}D^2=1.208 \pi cm^2

\Delta A=\frac{\pi}{4}(2.196cm)^2-1.2089 \pi cm^2=-0.00329\pi cm^2

Clearing Delta T

\Delta T=\frac{\Delta A}{2\alpha (A_o)}

\Delta T=\frac{-0.00329\pi cm^2}{2*1.2\times10^{-5}K^{-1}(1.208\pi cm^2)}=-113.47^0K

T_f=T_o-113.47^0K\Rightarrow T_f=-101.22^0C

/************************************

note: making the calculations by lineal dilatation the results isvery close

\Delta T=\frac{\Delta L}{\alpha (L_o)}=\frac{2.196-2.199}{1.2\times 10^{-5} (2.199)}=-113.68